Suppose the demand equation for computers by Teetan Ltd for the year 2017 is given
by Qd= 1200-P and the supply equation is given by Qs= 120+3P. Find equilibrium
price and analyse what would be the excess demand or supply if price changes to Rs
400 and Rs 120
1
Expert's answer
2020-12-18T10:41:00-0500

Teetan Ltd 2017

Price is at equilibrium when total demand equals the total supply.

But, total demand, Qd= 1200-P and total supply, Qs= 120+3P

Therefore, at Teetan Ltd’s equilibrium price, 1200-P = 120+3P

When putting like terms together, 1200- 120 = 3P + P

Hence, 1,080 = 4P

"P = \frac{1,080}{4} = 270"

At equilibrium, total demand = 1200-P = 1200 – 270 = Rs 930


Also at equilibrium, total supply = 120+3P = "(120) + (3 \times 270) = 120 + 810 = Rs 930"


The excess demand if price changes to Rs 400 and Rs 120

At Rs 400, total demand = 1200-P = 1200 – 400 = Rs 800

Since the demand is lower than the one at equilibrium, there will be no excess demand


At Rs 120, total demand = 1200-P = 1200 – 120 = Rs 1080.

Excess demand = the new demand less demand at equilibrium = Rs 1080 - Rs 930 = Rs 150

Demand will be excess by 150


The excess supply if price changes to Rs 400 and Rs 120

At Rs 400, total supply = 120+3P = "(120) + (3 \times 400) = Rs 1320"

The excess supply = the new supply less supply at equilibrium = Rs 1320 - Rs 930 = Rs 390

Supply will be excess by 390


At Rs 120, total supply = 120+3P = "(120) + (3 \times 120) = Rs 480"

Since the supply is lower than the one at equilibrium, there will be no excess supply